Jepoy has one 100-peso bill, two 200-peso bills, and five 500-peso bills in his wallet. He wants to randomly pick one bill. Let X be the random variable of choosing one bill from his wallet.



Questions:


1. Construct the probability distribution table of X. [2 points]


2. Compute for the mean of X, denoted as E(X). [2 points]


3. Interpret the mean of X. [1 point]


4. Compute for the variance of X, denoted as V(X). [5 points]


5. Compute for the standard deviation of X, denoted as SD(X). [2 points]


Interpret the standard deviation of X. [1 point

1
Expert's answer
2022-03-19T04:09:34-0400

"1:\\\\P\\left( X=100 \\right) =\\frac{1}{8}=0.125\\\\P\\left( X=200 \\right) =\\frac{2}{8}=0.25\\\\P\\left( X=500 \\right) =\\frac{5}{8}=0.625\\\\\\begin{matrix}\tx&\t\t100&\t\t200&\t\t500\\\\\tP\\left( X=x \\right)&\t\t0.125&\t\t0.25&\t\t0.625\\\\\\end{matrix}\\\\2:\\\\EX=100\\cdot 0.125+200\\cdot 0.25+500\\cdot 0.625=375\\\\3:\\\\At\\,\\,average\\,\\,375-peso\\,\\,bill\\,\\,is\\,\\,taken\\\\4:\\\\EX^2=100^2\\cdot 0.125+200^2\\cdot 0.25+500^2\\cdot 0.625=167500\\\\V\\left( X \\right) =EX^2-\\left( EX \\right) ^2=167500-375^2=26875\\\\5:\\\\SD\\left( X \\right) =\\sqrt{V\\left( X \\right)}=\\sqrt{26875}=163.936\\\\6:\\\\At\\,\\,average\\,\\,the\\,\\,range\\,\\,of\\,\\,bill\\,\\,amount\\,\\,is\\,\\,\\pm 164 peso"